Assume that a problem A cannot be solved in O(n2) time. However, we can transform A into a problem B in O(n2 log n) time, and then solve B, and finally transform the solution of B in O(n) time into a solution for A.

Prove or Disprove: The above approach shows that B cannot be solved asymptotically less than O(n2) time.
Answer #1

Given:

a) A cannot be solved in O(n2).

b) B can be solved in O(n2logn) and solution from B can be transformed in O(n) time.

Proof:

Total complexity of solution while solving problem using B is: Time complexity of Solution B + Time complexity of solution transformation from B

Time complexity of  solving problem using B = O(n2logn)+O(n) = O(n2logn+n)

Consider the omega notation of time complexity for lower bound

phpiWcEjb.png

(Source: Tutorialpoints)

Ω(f(n)) ≥ { g(n) : there exists c > 0 and n0 such that g(n) ≤ c.f(n) for all n > n0. }

Here f(n) is O(n2logn+n) and g(n) is O(n2logn) which satisfy the above condition at c>=1 and n0 = 0

n2logn < 1*(n2logn+n) for all n>0

So lower bound of solution B is g(n) i.e. O(n2logn)

So B cannot be solved asymptotically less than O(n2logn) time

Therefore B cannot be solved asymptotically less than O(n2) time